OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 30°C and 70°C. The temperature is at its lowest point when t = 0 and completes one cycle over a 10-hour period. What is a sine function that would model this reaction? @ganeshie8

OpenStudy (anonymous):

@lolabieber

OpenStudy (anonymous):

a. f(t) = 20 sin 10t + 50 b. f(t) = 50 sin 10t + 20 c. f(t) = 50 sin pi/5t + 20 d. f(t) = 20 sin pi/5t + 50

OpenStudy (anonymous):

@1lorax2

OpenStudy (anonymous):

I'm sorry, I don't know trig very well.